Md=25y-80r,Ms=2400, p=2



a)       Find equilibrium output equation. And Y* when r=12%



b)     Find LM curve equation



c)      Find slope of LM curve.



d)     MS increased to 3000



3. Find equilibrium r



4. Find equilibrium y




1
Expert's answer
2022-04-17T17:16:02-0400

a) "25\cdot y – 80\cdot 0.
12 = 2400"

y = 95.6


b) r = (1/L 2) [L 0 + L 1Y – M/P]


d) "25\cdot y – 80\cdot 0.12 = 3000"

y = 119.6


3) 0.12


4) y = 95.6